OpenStudy (anonymous):

helpp!! solve the equation √c=20. check your answer

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[\sqrt{c}=20\\ (\sqrt{c})^{2}=20^{2}\\c ^{(\frac{ 1 }{ 2 })^2}=20^{2}\\c^1=400\\ c=400 \]